# Stellar metallicity is a key parameter for the search of Life in the Universe

### v3

@@ 1 Introduction
Current data suggests that CHNOPS elements are present in a large number of organic chemical species in the interstellar medium, in the star-forming regions and in protoplanetary disks (Nomura et al., 2021), making the basic building blocks of life a non limiting factor in the emergence of Life in the cosmos. The energy needed for biomass synthesis is also abundant in the Universe. Stars are a long-lasting and abundant source of free energy sustaining dynamic planetary atmospheres and the eventual biospheres. For instance, free energy flux is not a limiting factor for production of biomass by means of (hypothetical) oxygenic photosynthesis on exoplanets around main-sequence stars hotter than 3000 K (Covone et al., 2021), and geological processes might sustain life through geochemical and thermal disequilibria without the need for light (Merino et al., 2019). Hence, the first two necessary ingredients for life appear to be abundant on the cosmic scene and potentially in all planetary systems. Also liquid water is likely not a limiting factor. Recent observational data and theoretical studies suggest that the presence of liquid water (including subsurface oceans), might be more abundant in the cosmos than previously estimated (Noack et al., 2017).
@@
This set of conditions includes a large number of possible astronomical targets. For instance, recent work predicts that the forthcoming PLATO mission (Rauer et al., 2024) will allow to discover $24\pm 4$ Earth-like exoplanets (i.e., planets with radius between 0.8 – 1.25 Earth radii) around FGK stars on long period orbits (between 250 and 500 days), roughly including the circumstellar habitable zone (Matuszewski et al., 2023). We remark that FGK stars offer much better conditions for Life in terms of stability (i.e., frequency of high-energy flares) and X-ray and UV irradiance with respect to the most numerous M star dwarfs. If we also consider larger exoplanets (i.e., sub-Neptunes), this figure increases enormously: $172\pm 15$ exoplanets with radii between 1.2 and 2 Earth radii (so called super-Earths) (Matuszewski et al., 2023). This set includes Hycean worlds, that is, worlds hosting water oceans at their surface beneath an H2-rich atmosphere. Hycean worlds broaden the generally considered range of habitability, in terms of planetary mass, radius and orbital distance (see for instance Rigby & Madhusudhan (2024)). The larger sizes of sub-Neptunes makes them favourable targets for atmospheric characterization and biosignatures search via transit spectroscopy. About one third of these systems could be around stars bright enough to allow follow-up for mass measurements (via Doppler spectroscopy) and atmospheric characterization (with the JWST), with larger exoplanets being favoured in this metric. Thus in a few years, discoveries from PLATO and the on-going TESS missions will likely exceed our current observational follow-up capacity. Prioritising promising targets using present-day knowledge of Earth’s life appears necessary for an effective astronomical search for biosignatures.
@@ 2 The role of stellar metallicity in habitability
Star metallicity $Z$ (i.e., the abundance by mass of metals in a star relative to the total mass of the star) appears to be correlated to the presence of giant planets (Brewer et al., 2018), probably because the prestellar enrichment of heavier elements in the interstellar matter is necessary for the formation of the giant planets’ rocky cores (see Fischer & Valenti (2005) and references therein). Statistical analyses of known exoplanets show no trend between rocky planets occurrence and host-star metallicity (Zhu & Dong, 2021), albeit this might change in the future as more exoplanets are discovered. Indeed, a too low metal content in the protoplanetary disk is expected to inhibit the fast accretion of rocky bodies massive enough to start aggregating the gas in the circumstellar disk before this disappears in a few million years. The lower metallicity threshold necessary for rocky planet formation is still unknown. Theoretical models suggest that the first Earth-like planets might have formed from protoplanetary disks with metallicities $Z\gtrsim 0.1Z_{\odot}$ (Johnson & Li, 2012).
@@ 3 Metal availability as a key control on Life emergence and evolution
These enzymes are evolutionarily tuned to be able to transfer electrons to and from target molecules in a controlled way (Falkowski et al., 2008), allowing life to take advantage of the energy associated with the reaction and to perform chemical (and often physical) work (Appendix B). Since redox reactions require the flow of electrons, oxidoreductases reaction centres need to be finely tuned in their midpoint electron potential to accept and donate electrons without dissipating unnecessary energy. The vast majority of known oxidoreductases uses a diverse set of metal containing cofactors to achieve this (Giovannelli, 2023). For example, oxidoreductases utilised to extract high energy electrons from hydrogen (a common electron donor in microbial redox reactions) contain generally a nickel-iron metal cofactor (see Fig. 2), while many enzymes acting on oxygen (a strong electron acceptor respired by a large number of organisms) use instead copper containing cofactors to be able to gently pass electron to the strongly electronegative O2 molecule (32). While most known enzymes use one or more first or second row transition metals (see Figure 1), the list of biologically important metal cofactors is growing (Giovannelli, 2023). Only recently we have identified enzymes that catalyse the first step of methane oxidation (an important biosignature and greenhouse gas) using the rare earth elements lanthanum and cerium. The exact composition of the metal cofactors used by life is still unknown, however the majority of known oxidoreductases involved in key energy conserving reactions contain transition metals that include Fe, Cu, Mo, V, W, Co, Ni, Mg, Mn among others (see Fig. 1) (Hay Mele et al., 2023). These biometals (sensu (Giovannelli, 2023)) are de facto the key component of the engines that life uses to tap into the thermodynamic disequilibria needed to fuel metabolism.
@@ 4 Stellar metallicity as a key parameter in the search for Life
While we search our solar system and beyond for targets of astrobiological interest, we argue that the availability of transition metals necessary to access a diverse array of thermodynamic disequilibria is an essential feature of habitability. Chemical elements heavier than helium (including all biometals) have been produced in detectable amounts relatively late in cosmic history (Johnson et al., 2020) (Fig. 1). Hydrogen and helium have been produced in the early primordial nucleosynthesis, while synthesis of heavier chemical elements could be possible in significant quantities only during several stages of stellar evolution, with the first generation of stars (the so-called Population III) appearing on the cosmic scene when the Universe was not younger than about 100-200 million years old (Omukai & Nishi, 1998). While details are still highly debated, this scenario is strongly supported by the measured chemical abundances in the oldest stars in the Galaxy and the molecular clouds (Nomoto et al., 2013).
All biometals lighter than iron are the outcomes of fusion nuclear reactions in the core of massive stars (i.e., with a mass larger than about two solar masses) (Johnson, 2019). As iron is the most tightly bound nucleus (with binding energy 8.8 MeV per nucleon), it marks a separation between lighter elements that can be synthesised by the fusion of nuclei and heavier elements that are produced via nuclear fission processes. For instance, most of the heavy elements up to bismuth are produced in this way in Type-Ia supernovae explosions (Nomoto et al., 2013). Supernovae explosions have also a key role in diffusing the nuclear ashes in the interstellar medium enriching molecular clouds.
@@
Measurements of stellar metallicity are increasingly available for elements heavier than iron, allowing for the investigation of metal distribution in the host stars of distant planetary systems. In particular, the Apache Point Observatory Galactic Evolution Experiment (APOGEE, Weinberg et al. (2019)) has completed an homogeneous, high-resolution and high signal-to-noise spectroscopic survey of the stellar populations of the Milky Way, allowing to obtain a detailed chemical composition for about 146,000 stars (see Figure 3). While the bulk planetary composition is controlled by the metal composition of the parent star (Adibekyan et al., 2021), the crustal abundance of metals on a planetary scale is controlled by a large number of local factors. Processes like core formation, crustal differentiation, active geology and redox conditions of the planetary surface all influence the distribution and availability of CHNOPS elements and metals (Wade et al., 2021), controlling their ultimate availability for Life. Despite this, first order estimates of bulk planetary composition might be enough to determine a first order approximation of the metal availability for life.
@@
A major issue in the search for life by means of astronomical observations is the quantification of the probability (i.e., the Bayesian likelihood) of a positive interpretation of the collected data. Data from a variety of sources need to be considered, given the ambiguous interpretation of any single biosignature (Catling et al., 2018). The likelihood of the collected data occurring in the presence of life should be weighted by including observed abundance or upper limits on the presence of biometals from the host star. While the knowledge of the exoplanet parameters (mass, insolation, radius, etc) and its atmosphere and climate strongly informs the evaluation of the posterior Bayesian probability for a positive detection of life, we pose that information about biometals availability should be considered.
@@ 5 Conclusions
In the near future, the ESA mission PLATO (Rauer et al., 2024) will find hundreds of candidate rocky planets in the temperate zone of Sun-like stars. Companion follow-up spectroscopic observations will confirm the exoplanetary interpretation of the transit signals and to determine their mass and bulk density (via Doppler shift technique). High-resolution (about a few tens of thousands) spectroscopic follow-ups aimed at measuring the presence of CHNOPS and biometals in the host-stars will be a valuable addition. This also includes planned follow-up observations for measuring radial velocities.
@@
Figure 1: Periodic table of elements and their nucleosynthesis. Biological utilisation of the elements and biometals (in orange and yellow) and their nucleosynthesis processes in the Universe. CHNOPS elements and elements used as electron donors/acceptors by biology are also reported. Iron and Manganese are the only biometals that not only serve as cofactors, but might be used as metabolic substrates for redox reactions.
@@
Figure 2: Example of metal containing enzymes involved in the production or utilisation of possible biosignature gases such as CH4, N2O and O2 . The three dimensional structure of the protein is showed as a continuous molecular surface (with significant transparency applied for visualisation purpose) coloured according to the different amino acid chains, while the cofactors are shown using a ball and stick model within the enzyme and on their own coloured according to the type of atoms. a, The key enzyme for methanogenesis, the Methyl-coenzyme M reductase from the archaea Methanopyrus kandleri (McrA, Protein Data Bank accession number 1E6V. Structures can be visualised and inspected at https://www.rcsb.org/). The nickel containing F430 cofactors are visible within the enzyme in two copies, together with other small ligands. The structure of the nickel (Ni2+) containing F430-cofactor (the nickel atom is coloured in green) responsible for the last reduction step required for methane production is visible below the enzyme. b, The enzyme responsible for the production of nitrous oxide, Nitrous oxide reductase from the bacterium Pseudomonas stutzeri (NosZ, PDB 3SBR) with the two copper containing cofactor visible within the enzyme. The structure of the copper sulphide cofactor (Cu4S2) and the dinuclear copper is visible below the enzyme. Copper atoms are coloured in orange while sulphur atoms are in yellow. c, The complex multimeric structure of the Photosystem II from the cyanobacterium Thermosynechococcus elongatus (PSII, PDB 1S5L), where the oxygen evolving complex responsible for the splitting of water in oxygenic photosynthesis is located. All the chlorophyll and pigments have been removed for visualisation purposes. The structure of the oxygen evolving complex CaMn4O4 cofactor, the site of water splitting and oxygen generation is visible below the enzyme. Calcium is coloured in green, while manganese and oxygen are coloured in purple and red respectively.
@@
Figure 3: Examples of the distribution of key biometals in the APOGEE star catalogue. The two panels show the abundances of iron, nickel, and carbon for planet-hosting and non-planet hosting stars. Stars with known planets (coloured in orange) span a smaller range in chemical abundances with respect to the general stellar population (grey points). The Sun is coloured blue. a, Relative abundance of nickel as a function of stellar iron-to-hydrogen abundance, usually used as a proxy for stellar metallicity. b, Relative abundance of nickel as a function of the star carbon-to-iron abundance. So far, astronomers focus on metallicity distribution to investigate the connection with formation of planets. Abundance data are essential to inform the selection and interpretation of the most promising targets for search for life.
@@
CH4 Methane, an important greenhouse gas and potentially a biosignature if present with strong oxidants (e.g., oxygen) Methanogenesis (production) Cu, Ni, Fe, W
Methanotrophy (consumption) Fe, Cu, Mo
O2 Oxygen, a strong oxidant produced on Earth through oxygenic photosynthesis. O3 can be used as an indicator of oxygen presence Oxygenic photosynthesis (production) Mn, Fe, Cu
Aerobic respiration (consumption) Cu, Fe
N2O Nitrous oxide, possible biosignature if present together with reductants (CH4, H2) in the atmosphere Denitrification, nitrous oxide reduction (consumption) Fe, Cu
NH3 Ammonia, possible biosignature in the atmosphere since detectable levels have to be sustained by a potentially biological surface production Denitrification to ammonia, nitrogen fixation (production) Fe, Mo, Fe, V
Ammonia oxidation (consumption) Fe, Cu
(CH3)2S, CH3Cl, CH3 SH Methyl containing gases, produced on Earth through biological pathways might be established as alternative biosignatures Carbon fixation (production) Fe, V

### v4

# Transition metal abundance as a key parameter for the search of Life in the Universe

@@ 1 Introduction
Current data suggests that CHNOPS elements are present in a large number of organic chemical species in the interstellar medium, in the star-forming regions and in protoplanetary disks (Nomura et al., 2021), making the basic building blocks of life a non limiting factor in the emergence of Life in the cosmos. The energy needed for biomass synthesis is also abundant in the Universe. Stars are a long-lasting and abundant source of free energy sustaining dynamic planetary atmospheres and the eventual biospheres. For instance, free energy flux is not a limiting factor for production of biomass by means of (hypothetical) oxygenic photosynthesis on exoplanets around main-sequence stars hotter than about 3000 K (see for instance Covone et al., 2021), and geological processes might sustain life through geochemical and thermal disequilibria without the need for light (Merino et al., 2019). Hence, the first two necessary ingredients for life appear to be abundant on the cosmic scene and potentially in all planetary systems. Also liquid water is likely not a limiting factor. Recent observational data and theoretical studies suggest that the presence of liquid water (including subsurface oceans), might be more abundant in the cosmos than previously estimated (e.g., Noack et al., 2017).
@@
This set of conditions includes a large number of possible astronomical targets. For instance, recent work predicts that the forthcoming PLATO mission (Rauer et al., 2025) will allow to discover $24\pm 4$ Earth-like exoplanets (i.e., planets with radius between 0.8 – 1.25 Earth radii) around FGK stars on long period orbits (between 250 and 500 days), roughly including the circumstellar habitable zone (Matuszewski et al., 2023). We remark that FGK stars offer much better conditions for Life in terms of stability (i.e., frequency of high-energy flares) and X-ray and UV irradiance with respect to the most numerous M star dwarfs. If we also consider larger exoplanets (i.e., sub-Neptunes), this figure increases enormously: $172\pm 15$ exoplanets with radii between 1.2 and 2 Earth radii (Matuszewski et al., 2023). This set includes Hycean worlds, that is, worlds hosting water oceans at their surface beneath an H2-rich atmosphere. Hycean worlds broaden the generally considered range of habitability, in terms of planetary mass, radius and orbital distance (see for instance Rigby & Madhusudhan (2024)). The larger sizes of sub-Neptunes makes them favourable targets for atmospheric characterization and biosignatures search via transit spectroscopy. About one third of these systems could be around stars bright enough to allow follow-up for mass measurements (via Doppler spectroscopy) and atmospheric characterization (with the JWST), with larger exoplanets being favoured in this metric. Thus in a few years, discoveries from PLATO and the on-going TESS missions will likely exceed our current observational follow-up capacity. Prioritising promising targets using present-day knowledge of Earth’s life appears necessary for an effective astronomical search for biosignatures.
@@ 2 The role of stellar metallicity in habitability
Star metallicity $Z$ (i.e., the abundance by mass of chemical elements heavier than helium relative to the total mass of the star) appears to be correlated to the presence of giant planets (Brewer et al., 2018), probably because the prestellar enrichment of heavier elements in the interstellar matter is necessary for the formation of the giant planets’ rocky cores (see Fischer & Valenti (2005) and references therein). Statistical analyses of known exoplanets show no trend between rocky planets occurrence and host-star metallicity (Zhu & Dong, 2021), albeit this might change in the future as more exoplanets are discovered. Indeed, a too low metal content in the protoplanetary disk is expected to inhibit the fast accretion of rocky bodies massive enough to start aggregating the gas in the circumstellar disk before this disappears in a few million years. The lower metallicity threshold necessary for rocky planet formation is still unknown. Theoretical models suggest that the first Earth-like planets might have formed from protoplanetary disks with metallicities $Z\gtrsim 0.1Z_{\odot}$ (Johnson & Li, 2012).
@@ 3 Metal availability as a key control on Life emergence and evolution
These enzymes are evolutionarily tuned to be able to transfer electrons to and from target molecules in a controlled way (Falkowski et al., 2008), allowing life to take advantage of the energy associated with the reaction and to perform chemical (and often physical) work (Appendix B). Since redox reactions require the flow of electrons, oxidoreductases reaction centres need to be finely tuned in their midpoint electron potential to accept and donate electrons without dissipating unnecessary energy. The vast majority of known oxidoreductases uses a diverse set of metal containing cofactors to achieve this (Giovannelli, 2023). For example, oxidoreductases utilised to extract high energy electrons from hydrogen (a common electron donor in microbial redox reactions) contain generally a nickel-iron metal cofactor (see Fig. 2), while many enzymes acting on oxygen (a strong electron acceptor respired by a large number of organisms) use instead copper containing cofactors to be able to gently pass electron to the strongly electronegative O2 molecule. While most known enzymes use one or more first or second row transition metals (see Fig. 1), the list of biologically important metal cofactors is growing (Giovannelli, 2023). Only recently we have identified enzymes that catalyse the first step of methane oxidation (an important biosignature and greenhouse gas) using the rare earth elements lanthanum and cerium. The exact composition of the metal cofactors used by life is still unknown, however the majority of known oxidoreductases involved in key energy conserving reactions contain transition metals that include Fe, Cu, Mo, V, W, Co, Ni, Mg, Mn among others, see Fig. 1 (Hay Mele et al., 2023). These biometals (sensu Giovannelli, 2023) are de facto the key component of the engines that life uses to tap into the thermodynamic disequilibria needed to fuel metabolism.
@@ 4 Stellar metallicity as a key parameter in the search for Life
While we search our solar system and beyond for targets of astrobiological interest, we argue that the availability of transition metals necessary to access a diverse array of thermodynamic disequilibria is an essential feature of habitability. Chemical elements heavier than helium (including all biometals) have been produced in detectable amounts relatively late in cosmic history (Johnson et al., 2020), see Fig. 1. Hydrogen and helium have been produced in the early primordial nucleosynthesis, while synthesis of heavier chemical elements could be possible in significant quantities only during several stages of stellar evolution, with the first generation of stars (the so-called Population III) appeared on the cosmic scene when the Universe was not younger than about 100-200 million years old (Omukai & Nishi, 1998). While details are still highly debated, this scenario is strongly supported by the measured chemical abundances in the oldest stars in the Galaxy and the molecular clouds (Nomoto et al., 2013).
All biometals lighter than iron are the outcomes of fusion nuclear reactions in the core of massive stars (i.e., with a mass larger than about two solar masses) see e.g. Johnson (2019). As iron is the most tightly bound nucleus (with binding energy 8.8 MeV per nucleon), it marks a separation between lighter elements that can be synthesised by the fusion of nuclei and heavier elements that are produced via nuclear fission processes. For instance, most of the heavy elements up to bismuth are produced in this way in Type-Ia supernovae explosions (Nomoto et al., 2013). Supernovae explosions have also a key role in diffusing the nuclear ashes in the interstellar medium enriching molecular clouds.
@@
Measurements of stellar chemical abundances are increasingly available for elements heavier than iron, allowing for the investigation of metal distribution in the host stars of distant planetary systems. In particular, the Apache Point Observatory Galactic Evolution Experiment (APOGEE, Weinberg et al. (2019)) has completed an homogeneous, high-resolution and high signal-to-noise spectroscopic survey of the stellar populations of the Milky Way, allowing to obtain a detailed chemical composition for about 146,000 stars (see Figure 3). While the bulk planetary composition is controlled by the metal composition of the parent star (Adibekyan et al., 2021), the crustal abundance of metals on a planetary scale is controlled by a large number of local factors. Processes like core formation, crustal differentiation, active geology and redox conditions of the planetary surface all influence the distribution and availability of CHNOPS elements and metals (Wade et al., 2021), controlling their ultimate availability for Life.
More in details, Wang et al. (2019) found that elements with low condensation temperatures (lower than Earth’s differentiation temperature) are depleted in the terrestrial environment compared to solar abundances. In contrast, elements with higher condensation temperatures match solar abundances.
Despite this, first order estimates of bulk planetary composition might be enough to determine a first order approximation of the metal availability for life.
@@
A major issue in the search for life by means of astronomical observations is the quantification of the probability (i.e., the Bayesian likelihood) of a positive interpretation of the collected data. Data from a variety of sources need to be considered, given the ambiguous interpretation of any single biosignature (Catling et al., 2018). The likelihood of the collected data occurring in the presence of life should be weighted by including observed abundance or upper limits on the presence of biometals from the host star. While the knowledge of the exoplanet parameters (mass, insolation, radius, etc) and its atmosphere and climate strongly informs the evaluation of the posterior Bayesian probability for a positive detection of life, we propose that information about biometals availability should be considered.
@@ 5 Conclusions
In the near future, the ESA mission PLATO (Rauer et al., 2025) will find hundreds of candidate rocky planets in the temperate zone of Sun-like stars. Companion follow-up spectroscopic observations will confirm the exoplanetary interpretation of the transit signals and to determine their mass and bulk density (via Doppler shift technique). High-resolution (about a few tens of thousands) spectroscopic follow-ups aimed at measuring the presence of CHNOPS and biometals in the host-stars will be a valuable addition. This also includes planned follow-up observations for measuring radial velocities.
@@
Figure 1: Periodic table of elements and their nucleosynthesis. Biological utilisation of the elements and biometals (in orange and yellow) and their nucleosynthesis processes in the Universe are indicated. CHNOPS elements and elements used as electron donors/acceptors by biology are also reported. Iron and Manganese are the only biometals that not only serve as cofactors, but might also be used as metabolic substrates for redox reactions.
@@
Figure 2: Example of metal containing enzymes involved in the production or utilisation of possible biosignature gases such as CH4, N2O and O2 . The three dimensional structure of the protein is showed as a continuous molecular surface (with significant transparency applied for visualisation purpose) coloured according to the different amino acid chains, while the cofactors are shown using a ball and stick model within the enzyme and on their own coloured according to the type of atoms. (a), The key enzyme for methanogenesis, the Methyl-coenzyme M reductase from the archaea Methanopyrus kandleri (McrA, Protein Data Bank accession number 1E6V. Structures can be visualised and inspected at https://www.rcsb.org/). The nickel containing F430 cofactors are visible within the enzyme in two copies, together with other small ligands. The structure of the nickel (Ni2+) containing F430-cofactor (the nickel atom is coloured in green) responsible for the last reduction step required for methane production is visible below the enzyme. (b), The enzyme responsible for the production of nitrous oxide, nitrous oxide reductase from the bacterium Pseudomonas stutzeri (NosZ, PDB 3SBR) with the two copper containing cofactor visible within the enzyme. The structure of the copper sulphide cofactor (Cu4S2) and the dinuclear copper is visible below the enzyme. Copper atoms are coloured in orange while sulphur atoms are in yellow. (c), The complex multimeric structure of the Photosystem II from the cyanobacterium Thermosynechococcus elongatus (PSII, PDB 1S5L), where the oxygen evolving complex responsible for the splitting of water in oxygenic photosynthesis is located. All the chlorophyll and pigments have been removed for visualisation purposes. The structure of the oxygen evolving complex CaMn4O4 cofactor, the site of water splitting and oxygen generation is visible below the enzyme. Calcium is coloured in green, while manganese and oxygen are coloured in purple and red respectively.
@@
Figure 3: Examples of the distribution of key biometals in the APOGEE star catalogue. The two panels show the abundances of iron, nickel, and carbon for planet-hosting and non-planet hosting stars. Stars with known planets (coloured in orange) span a smaller range in chemical abundances with respect to the general stellar population (grey points). This is likely a consequence of the correlation between stellar metallicity and presence of giant planets. The Sun is coloured blue. Panel a: Relative abundance of nickel as a function of stellar iron-to-hydrogen abundance, sometimes used as a proxy for stellar metallicity. Panel b:, Relative abundance of nickel as a function of the star carbon-to-iron abundance. So far, astronomers focus on metallicity distribution to investigate the connection with formation of planets. Abundance data are essential to inform the selection and interpretation of the most promising targets for search for life.
@@
CH4 Methane, an important greenhouse gas and potentially a biosignature if present with strong oxidants (e.g., oxygen) Methanogenesis (production) Ni, Fe, Co, W, Mo, Cu
Methanotrophy (consumption) Fe, Cu, Mo, Ni, W
O2 Oxygen, a strong oxidant produced on Earth through oxygenic photosynthesis. O3 can be used as an indicator of oxygen presence Oxygenic photosynthesis (production) Mn, Fe, Cu, Mg
Aerobic respiration (consumption) Fe, Cu, Mn, Zn
N2O Nitrous oxide, possible biosignature if present together with reductants (CH4, H2) in the atmosphere Denitrification, nitrous oxide reduction (consumption) Fe, Cu, Mn, Ni
NH3 Ammonia, possible biosignature in the atmosphere since detectable levels have to be sustained by a potentially biological surface production Denitrification to ammonia, nitrogen fixation (production) Mo, Fe, V
Ammonia oxidation (consumption) Cu, Fe, Mo, Ni
(CH3)2S, CH3Cl, CH3 SH Methyl containing gases, produced on Earth through biological pathways might be established as alternative biosignatures Carbon fixation (production), secondary metabolite production, halogenation Fe, Mo, V, C, Cu